what are the coordinates  of orthocenter?

Grade:11

1 Answers

Latika Leekha
9 years ago
Hello student,
Let ABC be the given triangle and AD, BE and CF be the three altitudes.
Now, to calculate the orthocentre,
Step 1: Firstly, find the slopes of sides AB, BC and CA using the formula
\frac{y_{2}-y_{1}}{x_{2}-x_{1}}
Step 2: Calculate the slopes of altitudes AD, BE and CF which are perpendicular to BC, CA and AB respectively.
Slope of altitude = -1/ slope of opposite side
Step 3: By suing the point slope form, calculate the equation for the altitudes with their respective coordinates. The point slope formula is given as y – y1 = m (x – x1)
Step 4: Finally, by solving the equations of any two altitudes, we can get the orthocentre of the triangle.